media%2F538%2F53881865-1ed4-44ab-9236-97
Answer #1

7. c. gram

8. e. 118

9. d. A heterogeneous mixture

10. 100 + 273.15 = 373.15 K

11. one is in grams and the other is in kg . Take both the values in the same unit

convert 254 g into kg

254 \times 1000 = 0.254 g

Total amount of sulfur is 0.254 + 0.032 = 0.286 Kg

12. maybe -40 degree celcius not sure (sorry)

13. density = mass / volume since we want mass = density * volume

= 7.8754 * (25.4*25.4*25.4)

= 0.129 Kg mL

14.1 L = 1000 cm3

15. Burettes

16. dont know